# BECS: A Privacy-Preserving Computing Sharing Mechanism in 6G Computing Power Network

### v1

@@ Abstract
5G networks provide secure and reliable information transmission services for the Internet of Everything, thus paving the way for 6G networks, which is anticipated to be an AI-based network, supporting unprecedented intelligence across applications. Abundant computing resources will establish the 6G Computing Power Network (CPN) to facilitate ubiquitous intelligent services. In this article, we propose BECS, a computing sharing mechanism based on evolutionary algorithm and blockchain, designed to balance task offloading among user devices, edge devices, and cloud resources within 6G CPN, thereby enhancing the computing resource utilization. We model computing sharing as a multi-objective optimization problem, aiming to improve resource utilization while balancing other issues. To tackle this NP-hard problem, we devise a kernel distance-based dominance relation and incorporated it into the Non-dominated Sorting Genetic Algorithm III, significantly enhancing the diversity of the evolutionary population. In addition, we propose a pseudonym scheme based on zero-knowledge proof to protect the privacy of users participating in computing sharing. Finally, the security analysis and simulation results demonstrate that BECS can fully and effectively utilize all computing resources in 6G CPN, significantly improving the computing resource utilization while protecting user privacy.
@@ I Introduction
In the upcoming 6G era, it is anticipated that everything will be intelligently connected, thereby facilitating various data-intensive applications. By integrating communication networks deeply with various vertical industries, 6G will facilitate unprecedented applications such as Holographic Integrated Sensing and Communication (HISC), Artificial General Intelligence (AGI), and Digital Twins (DT) [1]. This indicates that AI will be the most crucial technology for constructing a comprehensively intelligent 6G networks, enabling network services to evolve dynamically and autonomously in response to demand. Fully coordinating end-edge-cloud computing devices to construct a Computing Power Network (CPN) will emerge as the leading paradigm for supporting ubiquitous intelligent services in 6G networks [2]. In essence, computing tasks will choose the most suitable offloading environment based on service demands, promoting a highly dynamic and hybrid computing space that further stimulates the sharing of computing resources within 6G networks.
CPN achieves dynamic allocation and efficient utilization of computing resources by integrating and scheduling computing devices throughout the network. Lu et al. [2] utilized deep reinforcement learning to find the optimal task transfer and resource allocation strategies in wireless CPN. Chen et al. [3] proposed an on-demand computing resource scheduling model to achieve efficient task offloading in edge CPN. However, as network services advance, resource allocation based on one or two objectives has become infeasible. Therefore, formulating computing resource optimization, an NP-hard task, as a multi-objective optimization problem (MOOP) is an effective approach. Peng et al. [4] formulated complex task offloading in the IIoT as a four-objective MOOP and developed a method to dynamically allocate computing resources based on the Non-dominated Sorting Genetic Algorithm III (NSGA-III). Gong et al. [5] employed the multi-objective evolutionary algorithm based on decomposition (MOEA/D) to optimize a three-objective edge task offloading problem, aiming to minimize delay and maximize rewards. The unprecedented services and complex architectures of 6G networks necessitate that CPN manage and optimize computing resources across multiple dimensions and continuously improve resource utilization efficiency [6]. Therefore, it is essential to change the current supply and demand dynamics of computing resources and to implement dynamic resource sharing.
Blockchain is envisioned as a promising approach to support distributed network management in 6G. Unlike traditional centralized approaches, blockchain enables trusted interactions among unfamiliar nodes through consensus mechanisms [7]. Xie et al. [8] exploited the immutability of blockchain to propose a resource trading mechanism based on sharding and directed acyclic graphs for large-scale 6G networks, thereby enhancing resource utilization efficiency. Dai et al. [9] developed a task-driven decentralized edge resource sharing scheme for 6G IoT, leveraging the distributed characteristics of blockchain. Wang et al. [10] proposed a provable secure blockchain-based federated learning framework for wireless CPN, aimed at accelerating the convergence of federated learning and enhancing the efficiency of wireless CPN. Meanwhile, the Federal Communications Commission views blockchain as a promising solution to lower the costs associated with radio spectrum management [6]. Consequently, blockchain is expected to fully realize its potential in the management and sharing of 6G resources.
@@
6G networks will integrate AI to fully merge the physical and digital worlds, necessitating enhanced security and privacy [11]. Xu et al. [12] utilized blockchain to provide adequate security and privacy protection for resource trading. Wang et al. [13] proposed a blockchain-enabled cyber-physical system integrating cloud and edge computing to achieve secure computing offloading. Samy et al. [14] introduced a blockchain-based framework for task offloading, ensuring security, integrity, and privacy in mobile edge computing. Although leveraging the characteristics of blockchain can provide preliminary security and privacy protection for edge and cloud computing allocation and trading, device-to-device computing sharing in 6G networks will require more comprehensive solutions to ensure the security and privacy between devices [15].
@@
Motivated by previous research, this article introduces a computing allocation and trading mechanism for 6G CPN, named BECS, which utilizes Blockchain and Evolutionary algorithms for efficient Computing resource Sharing. Furthermore, it provides a pseudonym scheme to protect users’ privacy. Specifically, the main contributions of this article are summarized as follows:
  • We formulate computing allocation as a MOOP with six objectives, employing an evolutionary algorithm to balance the interplay among these objectives, thus achieving optimal allocation schemes for 6G CPN.
  • We propose a novel evolutionary algorithm, NSGA-III-KDR, which improves the dominance relation of NSGA-III by using the kernel distance to enhance diversity in addressing the computing allocation MOOP.
The remainder of this article is organized as follows. Section II introduces the system model and formulates the computing sharing problem. Section III describes NSGA-III-KDR and its application in solving the computing allocation MOOP. Section IV presents the proposed pseudonym scheme and computing trading. Section V analyzes the security and computational complexity of the proposed scheme. Section VI presents simulation results. Section VII concludes the article.
II System Model and Problem Formulation
II-A System Model
@@
II-B Communication Model
The NOMA-based 6G networks, regarded the most promising future multi-access technology, is considered in BECS. It allows multiple users to transmit on the same frequency band, thereby enhancing spectrum resource utilization [18]. This model primarily addresses the communication latency in wireless communications, thus overlooking the latency in wired connections. Let $g_{u}$ denote the instantaneous channel gain from $U_{u}$ to the offloading computing device. Without loss of generality, the channel gains for all $U$ devices are ordered in descending sequence: $g_{1}<\cdots<g_{u}<\cdots<g_{U}$. Using successive interference cancellation and decoding for the user with the highest received power, the data transmission rate for $U_{u}$ during the computing offloading is calculated via Shannon’s theorem as follows:
@@ II-C Computing Model
II-D Service Model
1) Computing occupancy: As a direct factor of computing utilization, it can be quantified as the number of all devices participating in computing sharing while ensuring the provision of all network services. It is defined as follows:
@@ II-D Service Model
3) Load balancing: As a key metric for evaluating computing device workloads, load balancing aims to equalize each device’s load to the average. Concentrating multiple tasks on a single device diminishes computational efficiency and elevates energy consumption. The standard deviation of the tasks and the computing capabilities of the devices can reflect whether they are load balanced [22]. Therefore, the load balancing of $D_{d}$ can be calculated as
@@
4) Sharing revenue: A suitable revenue can encourage the participation of computing devices in resource sharing, thereby enhancing the overall resource utilization. The logarithmic utility function is used to quantify the sharing revenue of $D_{d}$, which can be calculated as
@@ II-E Premise Assumptions
III Computing Allocation Based on NSGA-III-KDR
@@
III-A MOOP of Computing Allocation
@@
III-B Kernel Distance-based Dominance Relation
@@
where $\theta_{X_{1}X_{2}}$ represents the acute angle between the two candidate solutions $X_{1}$ and $X_{2}$, which can be calculated as $\theta_{X_{1}X_{2}}=\operatorname{arccos}\bigl{(}F(X_{1}),F(X_{2})\bigr{)}$, and $\overline{\theta}$ denotes the size of the niche to which each candidate solution belongs.
As the number of objectives increases, neither the Euclidean distance nor the Mahalanobis distance can accurately reflect the crowding degree between individuals [26]. Therefore, NSGA-III-KDR utilizes the kernel distance from the point $X$ to the ideal point $P_{ide}$ to measure the similarity between them in handling high-dimensional problems [27]. It can be calculated as follows:
@@
III-C Encoding of Computing Resources
@@
3:        Evaluate objective functions by (11), (12), (13), (14) and (15);
@@ III-E Optimum Selection
By executing NSGA-III-KDR in BECS, CRP can set $\psi_{d}$ more reasonably by analyzing the optimal solution. Simultaneously, CRR can select the most appropriate CRP for task offloading based on the optimal solution. Consequently, both parties engaged in resource sharing achieve reliable, dynamic, and secure computing allocation and trading.
IV Secure Computing Trading with Privacy Protection
@@ IV-A System Initialization
IV-B Computing Devices Registration
@@ IV-C Pseudonym Generation
IV-D Pseudonymous Computing Trading
@@
IV-E Block Generation and Pseudonymous Update
@@
V Security Analysis and Discussion
@@ V-C Certificate Security
V-D Computational Complexity of NSGA-III-KDR
@@
VI Simulation Results and Analysis
@@
VI-A Simulation of Proposed NSGA-III-KDR
@@
VI-B Simulation of Proposed Computing Allocation Scheme
@@
As resource utilization improved, Fig. 8 to Fig. 8 illustrate the changes in the other five objectives considered by the computing allocation MOOP. Except for load balancing, improvements in all other objectives are observed with the optimization based on NSGA-III-KDR, attributed to enhanced resource utilization. The superior diversity of NSGA-III-KDR facilitates a more even distribution of computing resources, as particularly evidenced by the significant increase in privacy entropy. Although NSGA-III exhibits similar trends, it is less effective than NSGA-III-KDR. Notably, with NSGA-III-KDR and NSGA-II, increased resource utilization results in higher delays and sharing revenue, while energy consumption decreases as more tasks are offloaded to $\mathbb{U}$, leveraging all available resources within 6G PCN comprehensively. The optimization strategy of NSGA-II-SDR favors offloading tasks to $\mathbb{E}$ and $\mathbb{C}$, significantly reduces delay, but uniquely results in positive growth in load balancing among the four algorithms. The optimal solution of MOEA/D surpasses all other algorithms, achieving the highest resource utilization and substantial improvements in all objectives, except for load balancing, and the decline in load balancing suggests a rational allocation of computing resources. However, the overall data of MOEA/D is the worst, with a negative optimization of 12.99%. Furthermore, although its median or average increase in resource utilization is less than that of NSGA-III-KDR, the more substantial improvement in load balancing suggests a less favorable overall evolution. Since the selection of the optimal solution using TOPSIS is random, not every iteration of MOEA/D yields a solution that outperforms those of other algorithms.
@@
VI-C Simulation of Proposed Pseudonymous Scheme

### v2

# BECS: A Privacy-Preserving Computing Sharing Mechanism for 6G Computing Power Network

@@ Abstract
The evolution towards AI-based 6G networks, building upon the connectivity foundation of 5G, promises unprecedented intelligent applications. However, realizing this vision necessitates a significant evolution in computing to effectively leverage distributed computing resources. This demand drives the development of the 6G Computing Power Network (CPN), which aims to aggregate and orchestrate heterogeneous computing resources spanning user devices, edge nodes, and cloud infrastructures. In this article, we propose BECS, a computing sharing mechanism based on evolutionary algorithm and blockchain, designed to balance task offloading among computing resources within 6G CPN, thereby enhancing the computing resource utilization. We model computing sharing as a multi-objective optimization problem, aiming to improve resource utilization while addressing other trade-offs. To tackle this NP-hard problem, we devise a kernel distance-based dominance relation and incorporate it into the Non-dominated Sorting Genetic Algorithm III, significantly enhancing population diversity. In addition, we propose a pseudonym scheme based on zero-knowledge proofs to protect user privacy during computing sharing. Finally, security analysis and simulation results demonstrate that BECS can effectively utilize all computing resources in the 6G CPN, significantly improving resource utilization while preserving user privacy.
@@ I Introduction
In the upcoming 6G era, it is anticipated that everything will be intelligently connected, enabling a wide range of data-intensive applications. By deeply integrating communication networks with various vertical industries, 6G will enable unprecedented applications such as Holographic Integrated Sensing and Communication (HISC), Artificial General Intelligence (AGI), and Digital Twins (DT) [1]. This indicates that AI will be one of the most crucial technologies for constructing a comprehensively intelligent 6G networks, enabling network services to evolve dynamically and autonomously in response to demand. The coordination of end-edge-cloud computing devices to form a Computing Power Network (CPN) is expected to become a leading paradigm for supporting ubiquitous intelligent services in 6G networks [2]. In essence, the CPN envisions intelligent orchestration, where computing tasks are dynamically offloaded to the most suitable execution environment based on real-time service demands. This fosters a highly dynamic and hybrid computing environment that promotes the sharing of computing resources across 6G networks.
While Multi-access Edge Computing (MEC) in 5G networks paved the way for offloading computing tasks to the network edge and unleashing potential for the Internet of Things (IoT) [3], its limitations become apparent when facing the demands of the 6G era. MEC is typically characterized by localized deployments, limited resources, and siloed management, all of which hinder the integration and orchestration of ubiquitously distributed, heterogeneous computing resources across the network. In the 6G era, networks are expected to scale up to support trillions of connections. Meanwhile, intelligent and diversified applications will not only require computing resources that far exceed previous levels but will also urgently demand a ubiquitous computing paradigm. This paradigm should enable real-time coordination of network-wide computing resources, dynamic on-demand orchestration, and seamless cross-domain resource allocation [4]. As a result, 5G MEC is inadequate for meeting future computing demands in terms of both scale and operational paradigms. Therefore, 6G networks require the CPN to deeply integrate and intelligently orchestrate dispersed, heterogeneous, and geographically distributed computing resources across the entire network. Such integration promotes full resource utilization and enables efficient, large-scale computing sharing [5].
Driven by the unprecedented features and services of future 6G networks, communication networks are increasingly exploring distributed or multi-center management models [6]. Furthermore, the 6G CPN aims to comprehensively mobilize vast heterogeneous devices, edge nodes, and cloud computing resources across organizational boundaries. This cross-domain and heterogeneous environment, which lacks a central authority of trust, poses significant challenges for the access, scheduling, trading, and management of computing resources [7]. Traditional centralized computing resource management architectures struggle to adapt to such large-scale, dynamically evolving distributed systems. These systems are not only inefficient but also vulnerable to single points of failure. Therefore, it is necessary to explore novel distributed network management and governance mechanisms to ensure that entities within the CPN interact and collaborate in an autonomous, reliable, and efficient manner. Blockchain, with its inherent decentralization and its ability to enhance trust and transparency among multiple stakeholders, emerges as a key enabling technology for future 6G network management [8]. It offers secure, transparent, and traceable solutions for computing sharing, trusted transactions, access control, and other critical functionalities within the CPN.
Furthermore, the ubiquitous sharing and collaboration of computing resources exacerbate the risk of user privacy leakage. The 6G CPN involves a wide variety of computing devices. In particular, when user devices participate in computing sharing, information such as user identity, behavioral patterns, and interaction records may be collected and analyzed by attackers to infer sensitive data or perform identity tracking [9]. These potential threats may discourage users from actively engaging in computing sharing. Therefore, to achieve efficient computing sharing, it is essential to implement secure and effective privacy protection strategies to safeguard user rights and establish a trustworthy CPN environment.
To address these intertwined challenges of efficiency, trust, and privacy, this article introduces a computing allocation and trading mechanism for 6G CPN, named BECS, which utilizes Blockchain and Evolutionary algorithms for efficient Computing resource Sharing. Furthermore, it provides a pseudonym scheme to protect users’ privacy. Specifically, the main contributions of this article are summarized as follows:
  • We formulate computing allocation as a multi-objective optimization problem (MOOP) with six objectives, employing an evolutionary algorithm to balance the interplay among these objectives, thus achieving optimal allocation schemes for 6G CPN.
  • We propose a novel evolutionary algorithm, NSGA-III-KDR, which improves the dominance relation of Non-dominated Sorting Genetic Algorithm III (NSGA-III) by using the kernel distance to enhance diversity in addressing the computing allocation MOOP.
The remainder of this article is organized as follows. Section II reviews related work. Section III introduces the system model and formulates the computing sharing problem. Section IV describes NSGA-III-KDR and its application in solving the computing allocation MOOP. Section V presents the proposed pseudonym scheme and computing trading. Section VI analyzes the security and computational complexity of the proposed scheme. Section VII presents simulation results. Section VIII concludes the article.
II Related Works
In this section, we present computing allocation and sharing in CPN/end-edge-cloud, and elaborate blockchain for decentralized management and privacy preservation.
II-A Computing Allocation and Sharing in CPN/End-Edge-Cloud
The dynamic allocation and efficient utilization of computing resources are central to the concept of CPN. Addressing the inherent optimization challenges in managing these aspects has been a key research focus. Lu et al. [2] utilized deep reinforcement learning to find the optimal task transfer and computing allocation strategies in wireless CPN. Chen et al. [5] proposed an on-demand two-stage computing resource scheduling model to achieve efficient task offloading in edge CPN. However, as 6G services become increasingly diverse and demanding, optimizing for single or dual objectives often proves insufficient to capture the inherent trade-offs between factors like latency, energy consumption, cost, and resource utilization [10].
Therefore, formulating computing allocation as a MOOP has become a more effective approach for handling these conflicting objectives. Meanwhile, evolutionary algorithms are widely adopted due to their ability to identify a set of non-dominated solutions representing different trade-offs. Peng et al. [11] formulated complex task offloading in the IIoT as a four-objective MOOP and developed a method to dynamically allocate computing resources based on the NSGA-III. Gong et al. [12] employed the multi-objective evolutionary algorithm based on decomposition (MOEA/D) to optimize a three-objective edge task offloading problem, aiming to minimize delay and maximize rewards. While these approaches demonstrate the applicability of MOEAs, standard algorithms like NSGA-III and MOEA/D can face challenges when dealing with many-objective optimization, often struggling to maintain sufficient population diversity alongside convergence pressure [13]. This limitation can hinder the exploration of the full solution space, particularly for complex 6G CPN scenarios involving numerous performance dimensions and heterogeneous resources.
II-B Blockchain for Decentralized Management and Privacy Preservation
Building on the need for decentralized trust and management in 6G networks, blockchain technology has been actively explored as a key enabler [14]. Xie et al. [15] exploited the immutability of blockchain to propose a resource trading mechanism based on sharding and directed acyclic graphs for large-scale 6G networks, thereby enhancing resource utilization efficiency. Nguyen et al. [16] proposed a blockchain-based mobile edge-cloud computation offloading scheme, leveraging the distributed characteristics of blockchain to to provide secure and trusted computing services. Wang et al. [17] proposed a provable secure blockchain-based federated learning framework for wireless CPN, aimed at accelerating the convergence of federated learning and enhancing the efficiency of wireless CPN. These works highlight blockchain’s potential to automate, secure, and streamline interactions like resource discovery, access control, scheduling coordination, and payment settlement via smart contracts [18], thereby fostering a reliable environment for large-scale computing sharing.
@@ II-B Blockchain for Decentralized Management and Privacy Preservation
6G networks will integrate AI to fully merge the physical and digital worlds, necessitating enhanced security and privacy [19]. Nguyen et al. [20] utilized blockchain to provide adequate security for task offloading in mobile edge computing. Wang et al. [21] proposed a blockchain-enabled cyber-physical system integrating cloud and edge computing to achieve secure computing offloading. Samy et al. [22] introduced a blockchain-based framework for task offloading, ensuring security, integrity, and privacy in mobile edge computing. Although leveraging the characteristics of blockchain can provide preliminary security and privacy protection for edge and cloud computing allocation and trading, device-to-device computing sharing in 6G networks will require more comprehensive solutions to ensure the security and privacy between devices [23].
Distinct from the aforementioned works, BECS is designed as a holistic mechanism for the 6G CPN. It uniquely combines a blockchain architecture, advanced multi-objective optimization techniques, and cryptographic privacy protection to tackle the intertwined challenges of efficiency and security in computing sharing of 6G CPN.
III System Model and Problem Formulation
@@
III-A System Model
@@
III-B Communication Model
The NOMA-based 6G networks, regarded as the most promising future multi-access technology, is considered in BECS. It allows multiple users to transmit on the same frequency band, thereby enhancing spectrum resource utilization [26]. This model primarily addresses the communication latency in wireless communications, thus overlooking the latency in wired connections. Let $g_{u}$ denote the instantaneous channel gain from $U_{u}$ to the offloading computing device. Without loss of generality, the channel gains for all $U$ devices are ordered in descending sequence: $g_{1}<\cdots<g_{u}<\cdots<g_{U}$. Using successive interference cancellation and decoding for the user with the highest received power, the data transmission rate for $U_{u}$ during the computing offloading is calculated via Shannon’s theorem as follows:
@@ III-C Computing Model
III-D Service Model
1) Computing occupancy: As a direct indicator of computing resource utilization, it is quantified by the number of devices participating in computing sharing while maintaining the provision of all network services. It aims to enhance the overall breadth of resource engagement across the system, which can be defined as follows:
@@ III-D Service Model
3) Load balancing: As a key metric for evaluating computing device workloads, load balancing aims to equalize each device’s load to the average while ensuring an equitable and efficient distribution of workload based on device capabilities. Concentrating multiple tasks on a single device reduces computational efficiency and increases energy consumption. The standard deviation of tasks and the computing capabilities of devices can indicate whether they are load-balanced [30]. Therefore, the load balancing of $D_{d}$ can be calculated as
@@
4) Sharing revenue: A suitable revenue can encourage the participation of computing devices in computing sharing, thereby enhancing the overall resource utilization. The logarithmic utility function is used to quantify the sharing revenue of $D_{d}$, which can be calculated as
@@ III-E Premise Assumptions
IV Computing Allocation Based on NSGA-III-KDR
@@
IV-A MOOP of Computing Allocation
@@
IV-B Kernel Distance-based Dominance Relation
@@
where $\theta_{X_{1}X_{2}}$ represents the acute angle between the two candidate solutions $X_{1}$ and $X_{2}$, which can be calculated as $\theta_{X_{1}X_{2}}=\operatorname{arccos}\bigl{(}F(X_{1}),F(X_{2})\bigr{)}$, and $\overline{\theta}$ denotes the size of the niche to which each candidate solution belongs, and can be set to the $\left\lfloor\frac{|P|}{2}\right\rfloor$-th minimum element of
$\left\{\min_{q\in p\setminus\{p\}}\theta_{pq}|p\in P\right\},$ (28)
where $\theta_{pq}$ is the acute angle between $p$ and $q$ of any pair of candidate solutions.
The kernel distance is chosen because, as the number of objectives increases, neither the Euclidean distance nor the Mahalanobis distance can accurately reflect the crowding degree between individuals [33]. Therefore, NSGA-III-KDR utilizes the kernel distance from the point $X$ to the ideal point $P_{ide}$ to measure the similarity between them in handling high-dimensional problems [34]. It can be calculated as follows:
@@
IV-C Encoding of Computing Resources
@@
3:        Evaluate objective functions by (13), (14), (15), (16), (17) and (18);
@@ IV-E Optimum Selection
By executing NSGA-III-KDR in BECS, CRP can set $\psi_{d}$ more reasonably by analyzing the optimal solution. Simultaneously, CRR can select the most appropriate CRP for task offloading based on the optimal solution. Consequently, both parties engaged in computing sharing achieve reliable, dynamic, and secure computing allocation and trading.
V Secure Computing Trading with Privacy Protection
@@ V-A System Initialization
V-B Computing Devices Registration
@@ V-C Pseudonym Generation
V-D Pseudonymous Computing Trading
@@
V-E Block Generation and Pseudonymous Update
@@
VI Security Analysis and Discussion
@@ VI-C Certificate Security
VI-D Computational Complexity of NSGA-III-KDR
@@
VII Simulation Results and Analysis
@@
VII-A Simulation of Proposed NSGA-III-KDR
@@
VII-B Simulation of Proposed Computing Allocation Scheme
@@
As resource utilization improved, Fig. 8 to Fig. 8 illustrate the changes in the other five objectives considered by the computing allocation MOOP. Except for load balancing, improvements in all other objectives are observed with the optimization based on NSGA-III-KDR, attributed to enhanced resource utilization. The superior diversity of NSGA-III-KDR facilitates a more even distribution of computing resources, as particularly evidenced by the significant increase in privacy entropy. Although NSGA-III exhibits similar trends, it is less effective than NSGA-III-KDR. Notably, with NSGA-III-KDR and NSGA-III, increased resource utilization results in higher delays and sharing revenue, while energy consumption decreases as more tasks are offloaded to $\mathbb{U}$, leveraging all available resources within 6G CPN comprehensively. The optimization strategy of NSGA-II-SDR favors offloading tasks to $\mathbb{E}$ and $\mathbb{C}$, significantly reduces delay, but uniquely results in positive growth in load balancing among the four algorithms. The optimal solution of MOEA/D surpasses all other algorithms, achieving the highest resource utilization and substantial improvements in all objectives, except for load balancing, and the decline in load balancing suggests a rational allocation of computing resources. However, the overall data of MOEA/D is the worst, with a negative optimization of 12.99%. Furthermore, although its median or average increase in resource utilization is less than that of NSGA-III-KDR, the more substantial improvement in load balancing suggests a less favorable overall evolution. Since the selection of the optimal solution using TOPSIS is random, not every iteration of MOEA/D yields a solution that outperforms those of other algorithms.
@@
VII-C Simulation of Proposed Pseudonymous Scheme